# Building and aligning a 10-plane light converter

### v1

@@ Introduction
By going beyond a single plane of light modulation, multi-plane light conversion (MPLC) overcomes this limitation and enables full control over multiple spatial modes of light (?). To this end, a series of phase masks separated by free-space propagation are used, where any desired transformation is encoded by appropriately optimizing the different phase masks, and diffraction enables the mixing of spatially separated areas of the beam. Since its first demonstration over a decade ago, MPLC has been used for various applications with classical light (?, ?, ?). More recently, this technology has been adopted for quantum information processing, where the spatial modes of single (?) and entangled  (?) photons are manipulated for applications in quantum interference, computation (?), and communication experiments (?). A similar multi-plane approach has also been adopted in the context of diffractive neural networks (?).
In this tutorial, we detail how to build and align a 10-plane programmable light converter (fig.1a,b). While fixed MPLCs with pre-designed phase masks have been thoroughly explored and are even commercially available for certain applications (?), programmable MPLCs having a large number of planes have only been recently demonstrated (?). Here, we first discuss how to build a 10-plane programmable MPLC using a single commercial SLM and simple optical components. We then describe a procedure that uses the programmability of the MPLC to perform a precise, yet relatively simple, alignment of the positions of the MPLC phase masks with respect to the manipulated light. As MPLCs can be built using common optical components and a single SLM, and algorithms for calculating the desired phase masks are already publicly available (?), we hope our detailed guide will assist interested groups in quickly adopting this exciting technology.
@@
Before building and aligning an MPLC, one should conduct simulations to find the optimal number of planes, their desired locations on the SLM, the free-space propagation distance, and the number of pixels in each phase mask required for their specific application. Nevertheless, to be concrete, we discuss the build and alignment procedure with our specific parameters and design, as the principles are universal and should be helpful regardless of the particular parameters chosen.
@@
A prism is used to access the bottom part of the SLM after five reflections. The side of the prism is glued to a metal plate and mounted onto a kinematic mount (Thorlabs, KMSR). The height of the prism is controlled using a translation stage (Thorlabs, LX10). We place the prism so that no light impinges onto its center to avoid the $\approx 250\mu m$-wide region of poor optical quality. Due to mechanical constraints, the prism is placed at a distance of $69mm$ from the SLM, which is considered in calculating the phase masks. Still, we found using a prism superior to other retro-reflection solutions due to relatively minimal induced aberrations and polarization rotation. After the last five phase masks, at the output of the MPLC, a square 1/2” mirror is used to redirect the light toward the rest of the experimental setup.
Figure 1: 10-plane light converter. (a) 10 phase masks separated by free-space propagation are implemented using a single SLM, a mirror, and a prism. The light bounces between the SLM and the MPLC mirror five times on the top part and then five times on the bottom part. (b) Two pictures of the experimental implementation showing the SLM, prism, MPLC mirror, and output mirror. (c) A sketch of the first part of the experimental setup. In our experiment, a binary amplitude mask defines spatially separated modes that are imaged to the first plane of the MPLC (?). Two mirrors, one imaged onto the MPLC and one in the Fourier plane, are used to independently control the input angle and position of the light, respectively.
@@
After discussing the components and construction of the MPLC, we can now describe its alignment. While only three main elements are involved, the number of planes and the sensitivity of the MPLC to phase mask positioning errors, on the order of a few pixels, make precise alignment crucial and challenging. Luckily, as we will now discuss, the programmability of the SLM itself helps us to set a reliable and easy-to-implement procedure for aligning the MPLC. This procedure is general and could work for an arbitrary number of planes.
The alignment consists of three main stages: (a) rough alignment of the first five planes to get the light through the first part of the MPLC, (b) rough alignment of the last five planes to get the light to the output, and (c) fine alignment of the entire MPLC. At the end of this procedure, the centers of the light beam at the different planes and the phase mask on the SLM will coincide such that the transformations in the experiment should reproduce those designed in the simulation.
@@
After the light passes through the entire MPLC, finer alignment must be performed to find the exact position where each phase mask should be placed on the SLM. This is a crucial step, as one must ensure that the center of the beam hits the center of the simulated phase mask in each plane. For this, add a camera at the output of the MPLC and use appropriate imaging lenses so that it can be translated to directly image the last three planes (Fig. 2a). In our case, we calculate the MPLC transformations so that the output modes are located at plane ’eleven’ which is positioned $87mm$ after the last physical MPLC plane. Thus, we set the camera and imaging lenses to images planes 9, 10, and 11.
@@
The first step of the fine alignment is to find where the center of the beam hits the SLM on the first plane. Therefore, we need to be able to image the first plane of the MPLC to the camera. For this, we use the MPLC itself: add quadratic phase masks that implement Fresnel lenses on appropriate planes of the MPLC to image the first plane of the MPLC to planes 9-11, which the camera can directly image. For example, the first plane can be imaged onto the eleventh plane by adding lenses at planes 2, 4, and 8 (Fig.2a,b). The lenses at planes 2 and 4 implement a 4f system between planes 1 and 5, which is then imaged to plane 11 using the single lens at plane 8. At this point we do not yet know the exact distance between the SLM and the MPLC mirror, so the focal lengths of the lenses are only approximate. For our purpose at this point it is sufficient that they reasonably image the sharp features of the incident beam onto the camera.
@@
The alignment of the second plane is carried out in a similar manner by imaging it to the camera using other planes and using $\pi$-steps. The accuracy here will largely depend on the ability to precisely define the center of the beam in this plane. This highly depends on the specific modes used in the experiment. One can also try to add a phase mask on the first plane (for which we already know the exact center) to make the center of the beam at the second plane clearer. If the second plane is not in the desired location, tilt mirror M1 to correct for this, and verify that the position of the first plane has not changed.
@@
Before proceeding to find the positions of the last five planes, use the first five planes to estimate the spacing between the MPLC mirror and the SLM. For that, image the first plane to the fifth plane again using a 4f system with planes two and four. Add a $\pi$-step in both planes, and change the effective focal length of the lenses until reaching ideal imaging between the two planes, such that both $\pi$-steps are sharp. The focal length of the lenses corresponds to the distance traveled by the light between planes. In our system, we found this method gives an estimate with an accuracy of less than $1~{}mm$, where such an error did not affect the transformations we performed with the MPLC in a significant manner.
@@
Figure 2: Fine alignment of the MPLC. (a,b) The position of the first plane of the MPLC is found by imaging it onto the camera using effective lenses on planes 2, 4 and 8, and a $\pi$-step on plane 1. A correction pattern (not shown) is added to the SLM during the alignment process to compensate for its aberrations. (c) Single SLM-pixel accuracy is achieved by centering a $\pi$-step with the center of the beam.
@@
Finally, once the alignment is completed, we further correct for aberrations in the optical setup. Adding the SLM’s correction pattern (provided by Hamamatsu in our case) during the alignment improves its accuracy. We then further correct other aberrations after the alignment process by empirically adding spatially-varying (in our case, mode-dependent) phases at the first plane of the MPLC.

### v2

@@ Introduction
By going beyond a single plane of light modulation, multi-plane light conversion (MPLC) overcomes this limitation and enables full control over multiple spatial modes of light (?). To this end, a series of phase masks separated by free-space propagation are used, where any desired transformation is encoded by appropriately optimizing the different phase masks, and diffraction enables the mixing of spatially separated areas of the beam. Since its first demonstration over a decade ago, MPLC has been used for various applications with classical light (?, ?, ?, ?, ?, ?, ?). More recently, this technology has been adopted for quantum information processing, where the spatial modes of single (?) and entangled  (?) photons are manipulated for applications in quantum interference, computation (?), and communication experiments (?, ?). A similar multi-plane approach has also been adopted in the context of diffractive neural networks (?, ?, ?).
In this tutorial, we detail how to build and align a 10-plane programmable light converter (fig.1a,b). While fixed MPLCs with pre-designed phase masks have been thoroughly explored and are even commercially available for certain applications (?), programmable MPLCs having a large number of planes have only been recently demonstrated (?, ?). Here, we first discuss how to build a 10-plane programmable MPLC using a single commercial SLM and simple optical components. We then describe a procedure that uses the programmability of the MPLC to perform a precise, yet relatively simple, alignment of the positions of the MPLC phase masks with respect to the manipulated light. As MPLCs can be built using common optical components and a single SLM, and algorithms for calculating the desired phase masks are already publicly available (?), we hope our detailed guide will assist interested groups in quickly adopting this exciting technology.
@@
Before building and aligning an MPLC, one should conduct simulations to find the optimal number of planes, their desired locations on the SLM, the free-space propagation distance, and the number of pixels in each phase mask required for their specific application. To do so, one could start from an educated initial guess for the different parameters. For example, the distance between planes could be set such that the different modes start to diffract significantly (e.g. a few Rayleigh ranges for a Gaussian beam), the number of planes can be set to be larger than the number of modes if arbitrary transformations are desired, and the number of pixels per mode in each phase mask can be set to achieve decent sampling. Once a set of initial parameters have been determined and a target transformation has been identified, corresponding phase masks can be calculated via wavefront-matching  (?) or gradient descent  (?) methods. The fidelity of the transformation can then be optimized by tweaking the different parameters, choosing the ones to be implemented in the experiment.
To be concrete, we discuss the build and alignment procedure with our specific parameters and design, as the principles are universal and should be helpful regardless of the particular parameters chosen.
@@
A prism is used to access the bottom part of the SLM after five reflections. The side of the prism is glued to a metal plate and mounted onto a kinematic mount (Thorlabs, KMSR). The height of the prism is controlled using a translation stage (Thorlabs, LX10). We place the prism so that no light impinges onto its center to avoid the $\approx 250\mu m$-wide region of poor optical quality. Due to mechanical constraints, the prism is placed at a distance of $69mm$ from the SLM, which is considered when calculating the phase masks using wavefront-matching, by setting the distance between planes 5 and 6 to $2\cdot 69mm$. Still, we found using a prism superior to other retro-reflection solutions due to relatively minimal induced aberrations and polarization rotation. After the last five phase masks, at the output of the MPLC, a square 1/2” mirror is used to redirect the light toward the rest of the experimental setup.
Figure 1: 10-plane light converter. (a) 10 phase masks separated by free-space propagation are implemented using a single SLM, a mirror, and a prism. The light bounces between the SLM and the MPLC mirror five times on the top part and then five times on the bottom part. (b), (d) Side and top view pictures of the experimental implementation showing the SLM, prism, MPLC mirror, and output mirror. (c) A sketch of the first part of the experimental setup. M1 and M2 are mirrors, L1-L4 are lenses. In our experiment, a binary amplitude mask defines spatially separated modes that are imaged to the first plane of the MPLC (?). Two mirrors, one imaged onto the MPLC and one in the Fourier plane, are used to independently control the input angle and position of the light, respectively.
@@
After discussing the components and construction of the MPLC, we can now describe its alignment. While only three main elements are involved, the number of planes and the sensitivity of the MPLC to phase mask positioning errors, on the order of a few pixels, make precise alignment crucial and challenging. Luckily, as we will now discuss, the programmability of the SLM itself helps us to set a reliable and easy-to-implement procedure for aligning the MPLC  (?). This procedure is general and could work for an arbitrary number of planes.
The alignment consists of three main stages: (a) rough alignment of the first five planes to get the light through the first part of the MPLC, (b) rough alignment of the last five planes to get the light to the output, and (c) fine alignment of the entire MPLC. At the end of this procedure, the centers of the light beam at the different planes and the phase mask on the SLM will coincide such that the transformations in the experiment should reproduce those designed in the simulation. In our case, the light source used for the alignment and experiments with the MPLC are spatially entangled photons generated via Spontaneous Parametric Down-Conversion (SPDC)  (?, ?). Nonetheless, since the alignment relies on intensity measurements with a camera rather than correlations between single-photon detectors, the SPDC source effectively behaves as a spatially incoherent source. Indeed, the procedure below is general and can also be performed with classical light (ideally, linearly polarized for our polarization-sensitive liquid-crystal SLM  (?, ?)).
@@
After the light passes through the entire MPLC, finer alignment must be performed to find the exact position where each phase mask should be placed on the SLM. This is a crucial step, as one must ensure that the center of the beam hits the center of the simulated phase mask in each plane. For this, add a camera at the output of the MPLC and use appropriate imaging lenses so that it can be translated to directly image the last three planes (Fig. 2a). In our case, we calculate the MPLC transformations so that the output modes are located at plane ’eleven’ which is positioned $2\cdot 43.5mm$ after the last physical MPLC plane, such that the propagation distance between planes ten and eleven is equal to twice the SLM-mirror spacing. Thus, we set the camera and imaging lenses to image planes 9, 10, and 11.
@@
The first step of the fine alignment is to find where the center of the beam hits the SLM on the first plane. Therefore, we need to be able to image the first plane of the MPLC to the camera. For this, we use the MPLC itself: add quadratic phase masks that implement Fresnel lenses on appropriate planes of the MPLC to image the first plane of the MPLC to planes 9-11, which the camera can directly image. For example, the first plane can be imaged onto the eleventh plane by adding lenses at planes 2, 4, and 8 (Fig.2a,b). The lenses at planes 2 and 4, each with a focal length that is equal to the spacing between planes, implement a 4f system between planes 1 and 5, which is then imaged to plane 11 using the single lens at plane 8 with a focal length $f$ such that $\frac{1}{u}+\frac{1}{v}=\frac{1}{f}$ holds for distance $u$ to plane 5 and $v$ to plane 11. At this point we do not yet know the exact distance between the SLM and the MPLC mirror, so the focal lengths of the lenses are only approximate. For our purpose at this point it is sufficient that they reasonably image the sharp features of the incident beam onto the camera.
@@
The alignment of the second plane is carried out in a similar manner by imaging it to the camera using other planes and using $\pi$-steps. The accuracy here will largely depend on the ability to precisely define the center of the beam in this plane. This highly depends on the specific modes used in the experimental application, which we assume are the same ones used for alignment. If the center of the beam is hard to define in this plane, the already-aligned first plane of the MPLC can be used to add a phase mask (e.g., a Fresnel lens) to change the beam shape on the second plane and make its center easier to define. If the second plane is not in the desired location, tilt mirror M1 to correct for this, and verify that the position of the first plane has not changed.
@@
Before proceeding to find the positions of the last five planes, use the first five planes to estimate the spacing between the MPLC mirror and the SLM. For that, image the first plane to the fifth plane again using a 4f system with planes two and four. Add a $\pi$-step in both planes, and change the effective focal length of the lenses until reaching ideal imaging between the two planes, such that both $\pi$-steps are sharp. The focal length of the lenses corresponds to the distance traveled by the light between planes. In our system, we found this method gives an estimate with an accuracy of less than $1~{}mm$, where such an error did not affect the transformations we performed with the MPLC in a significant manner. The accuracy of this method was estimated by finding the focal length where clear defocus of the $\pi$-step is observed. This was then verified in experiments with phase masks calculated for different distances between planes, yielding the same optimal distance up to less than $1~{}mm$.
@@
Figure 2: Fine alignment of the MPLC. (a,b) The position of the first plane of the MPLC is found by imaging it onto the camera using effective lenses on planes 2, 4 and 8, and a $\pi$-step on plane 1. A correction pattern (not shown) is added to the SLM during the alignment process to compensate for its aberrations. (c) Single SLM-pixel accuracy is achieved by centering a $\pi$-step with the center of the beam. In our case, the light source consists of entangled photons passing through an array of 50 circular apertures, yielding the shape seen in the picture. The $\pi$-step on plane 1 is centered with the beam when it is centered with the middle column of light spots. On the right, upon moving the $\pi$-step by one pixel, clear asymmetry can already be seen on the spots at the center.
@@
Finally, once the alignment is completed, we further correct for aberrations in the optical setup. Adding the SLM’s correction pattern (provided by Hamamatsu in our case) during the alignment improves its accuracy. We then further correct other aberrations after the alignment process by empirically adding spatially-varying (in our case, mode-dependent) phases at the first plane of the MPLC. In addition, to find the optimal axial position of the detectors (in our case, single-photon detectors), which should be placed at plane eleven, we use a camera as well. We image once more a $\pi$-step from one of the MPLC planes to plane eleven using Fresnel lenses and find the position where the $\pi$-step is the sharpest using a camera. We place our detectors in this plane.
MPLC performance
The performance of a given MPLC depends on many factors, from the desired application, number of planes, and quality of alignment to pixel crosstalk and laser bandwidth. While some of these parameters are highly application-specific and can vary between experiments, we use numerical simulations to provide some intuition for the dependence of the performance on the number of planes and quality of alignment.
Preliminary studies into the required number of MPLC planes as a function of the number of spatial modes have shown a linear scaling (?), as one would expect from a naive counting of degrees of freedom. Furthermore, increasing the number of planes has been shown to increase the spectral bandwidth of the device and reduce the effect of pixel crosstalk (?). It is therefore generally advisable to increase the number of MPLC planes, as long as the device’s loss remains acceptable. We show a similar trend here by looking at the simulated fidelity of random unitary transformations on five Gaussian spatial modes as a function of the number of programmable phase planes (fig. 3a). While the specific values of the fidelity could change significantly when taking into account different SLM-mirror spacings, pixel crosstalk strengths, and other imperfections, the trend of improved performance with an increasing number of planes is representative. This motivated our choice of building a large MPLC that takes advantage of the entire size of our SLM.
Increasing the number of planes can, however, induce larger alignment errors, which can lower the fidelity of the programmed transformations. From our alignment procedure described above, we can identify two leading types of errors in the positioning of the phase masks. The first is simply a random error in the position of each mask, resulting from the finite accuracy of finding the center of the beam on each plane. As shown above, this error is at the single-pixel level in our case. Another error is a systematic one, resulting from a random error in the position of one plane propagating to the following ones. This type of error is not corrected by our alignment procedure, which uses previous planes for the alignment of a given plane. The extreme case is an error in the position of the second plane, which will linearly propagate to all following planes. For example, a one-pixel error in the second plane means that the third plane will be aligned using a Fresnel lens with an offset of one-pixel, causing a two-pixel error.
Thankfully, as one can see in Fig. 3b, errors propagating between planes, even when amounting to large errors on the last planes of the MPLC, do not significantly degrade the fidelity compared with random errors in each plane, motivating our choice of alignment method. We note that our approach is similar to the alignment of any optical setup. Each element is positioned as precisely as possible relative to the beam impinging on it, minimizing random errors in the position of each element. Errors can, however, propagate between optical elements. A shifted lens at the beginning of the setup can cause large beam deviations at the end of it, but the performance of the setup is not significantly degraded as long as all other components are aligned with respect to the shifted beam.
Figure 3: MPLC scaling and the effect of errors. (a) The fidelity of a $5\times 5$ random unitary transformation implemented using a different number of phase planes with equal spacings of $87mm$. The error bars represent the standard deviation for the fidelity from different random unitaries. (b) For 10 phase planes, the fidelity of realizing a 5-dimensional discrete Fourier transform is presented as a function of position errors (in the horizontal direction) of the center of the masks. For random position errors, the x-axis marks the maximal position error for each mask, assuming uniform distribution of errors up to this number. For propagated errors from the second plane, the x-axis is the actual error applied to the second plane, which is then linearly propagated to the following planes. The effect of random position errors in each plane is stronger than that of a propagated error from an error in the second plane. All fidelities are calculated according to the Frobenius norm and are presented after correcting mode-dependent phases at the input and output planes.